How do you convert z to x and x to z in Statistics? An explanation would be great. Thank you!
Answer #1

There are two formula we need to remember:

\small z=\frac{x-\mu}{\sigma}

and \small x=\mu+z*\sigma

We normally have three values given out of four \small x, z, \mu, \sigma and we need to find the 4th one.

\small \mu is mean, \small \sigma is population standard deviation.

We just need to use these and we can convert x to z or z to x as per our requirement.

Example:

mean = 5, population standard deviation = 2, x = 7 then z = ?

\small z=\frac{7-5}{2}=\frac{2}{2}=1

Same way, mean = 5, sd = 2, z = 1 then x = ?

\small x=5+1*2 = 5+2=7

That's how we convert. Please remember \small \sigma is population standard deviation.

I hope this will help. Thank you.